OpenStudy (kitkat1):

A paper popcorn cone measures 4 inches in diameter and is 9 inches high. What is the volume of popcorn that can be contained within this cone? Use 3.14 for pi. Round your answer to the nearest hundredth. 37.68 cubic inches 41.90 cubic inches 40.23 cubic inches 39.45 cubic inches

OpenStudy (kitkat1):

Please explain me how to do this Ill give medals

OpenStudy (kitkat1):

Anyone help asap

OpenStudy (kitkat1):

Can someone help me

OpenStudy (anonymous):

\[Volume = \pi r^2 (h/3)\] 4in diameter = 2in. radius h = 9

OpenStudy (kitkat1):

and then?

OpenStudy (anonymous):

put the value of r and h and solve use pi = 3.14

OpenStudy (kitkat1):

so V=1 (3.14)( 9)^2 (9) 3

OpenStudy (anonymous):

\[V = 3.14 *(2)^2 *(9/3) = 3.14 *4 *3\]

OpenStudy (kitkat1):

And what do I do now?

OpenStudy (anonymous):

multiply them

OpenStudy (kitkat1):

I got 37.68

OpenStudy (kitkat1):

Is tha right?

OpenStudy (kitkat1):

right no?

OpenStudy (anonymous):

correct

OpenStudy (kitkat1):

can u help me with one more?